OpenStudy (anonymous):

solve [-3.46-(-2)]^2-0.09

OpenStudy (anonymous):

Order of operations: Parenthesis, then exponents, then multiplication/division, then addition/subtraction.

OpenStudy (anonymous):

Only other trick is to remember that subtracting a negative number is like adding that number.

jhonyy9 (jhonyy9):

i think this is (-1,46)^2 -0,09 = 2,1316-0,09=2,0416